OpenStudy (yumi):

eqaution of line m=-1/2 (1,-2) can you plaese answer it guys?

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] Given gradient, \[m=-\frac{ 1 }{ 2 }\]and\[(1,-2)\] Thus, \[y-(-2)=(-\frac{ 1 }{ 2 } )(x-1)\]\[y+2=-\frac{ 1 }{ 2 }x +\frac{ 1 }{ 2 }\] \[y=-\frac{ 1 }{ 2 }x-\frac{ 3 }{ 2 }\]

OpenStudy (yumi):

thanks for helping me ^_^